A wheel rotates without friction about a stationary horizontal axis at the center of the wheel. A constant tangential force equal to 88.0 N is applied to the rim of the wheel. The wheel has radius 0.110 m . Starting from rest, the wheel has an angular speed of 13.9 rev/s after 3.28 s .

What is the moment of inertia of the wheel?

Express your answer with the appropriate units.
Answer #1

Torque is calculated as follows:

T = 88 N x 0.11 m = 9.68 Nm

Angular acceleration is calculated as follows:

dω/dt = (13.9 x 2π) /3.28 s = 26.6 rad/s²

moment of inertia is,

I = T / (dω/dt) = 9.68 / 26.6 = 0.364 kg.m²